OpenStudy (anonymous):

In the problem 3/4 • N = 1, N must be _____. A. the additive inverse of 3/4 B. the multiplicative inverse of 3/4 C. the additive identity D. the multiplicative identity

OpenStudy (automaticloveletter):

I'm going to go with B

OpenStudy (automaticloveletter):

The Multiplicative Inverse Property: The multiplicative inverse of a number, a is 1/a a x 1/a = 1 and 1/a x a = 1 This is often written in one line... a x 1/a = 1 = 1/a x a Where a is any real number.

OpenStudy (anonymous):

i don't know i was going to go will a and then i was thinking nah I'm going to ask the smart people

OpenStudy (anonymous):

thanks